A large tank contains 100 gallons of pure water. How many gallons of a saline solution containing 25% salt must be added to obtain a solution that has 10% salt ?
Answer by stanbon(75887)

A large tank contains 100 gallons of pure water. How many gallons of a saline solution containing 25% salt must be added to obtain a solution that has 10% salt ?
--------------------------------------
Equation:
salt + salt = salt
0*100 + 0.25x = 0.10(x+100)
Multiply thru by 100 and solve for "x"
0 + 25x = 10x + 1000
15x = 1000
x = 66 2/3 gallons of 25% saline solution must be add
=====================
Cheers,
Stan H.